Three events b1 b2 and b3 are mutually exclusive and exhaustive. If P(b1) = p(b2) = p(b3) and in the context of these events A is define such that p(A/B1) = 0.75 p(A/B2) = 0.93 p(A/B3)
Find p(b1/a ) and p(b3/a)
1
Expert's answer
2019-11-12T12:19:48-0500

Use Bayes' rule

Suppose "B_1,B_2,..., B_n" are mutually exclusive and exhaustive cases (hypotheses) and "A" is some event (evidence) that has been observed. Then


"P(B_i|A)={P(B_i)P(A|B_i) \over \displaystyle\sum_{j=1}^nP(B_j)P(A|B_j)}, 1\leq i\leq n"

"P(B_1|A)={P(B_1)P(A|B_1) \over P(B_1)P(A|B_1)+P(B_2)P(A|B_2)+P(B_3)P(A|B_3)}"

"P(B_2|A)={P(B_2)P(A|B_2) \over P(B_1)P(A|B_1)+P(B_2)P(A|B_2)+P(B_3)P(A|B_3)}"


"P(B_3|A)={P(B_3)P(A|B_3) \over P(B_1)P(A|B_1)+P(B_2)P(A|B_2)+P(B_3)P(A|B_3)}"

Given that


"P(B_1)=P(B_2)=P(B_3)""P(A|B_1)=0.75,P(A|B_2)=0.84,P(A|B_3)=0.93"

We assume that "P(B_1)=P(B_2)=P(B_3)\not=0". We don't need to compute "P(B_1),P(B_2),P(B_3)"



"P(B_1|A)={0.75\cdot P(B_1)\over 0.75\cdot P(B_1)+0.84\cdot P(B_2)+0.93\cdot P(B_3)}="

"={0.75\cdot P(B_1)\over 0.75\cdot P(B_1)+0.84\cdot P(B_1)+0.93\cdot P(B_1)}="

"={0.75\over 0.75+0.84+0.93}={25 \over 84}"



"P(B_2|A)={0.84\cdot P(B_2)\over 0.75\cdot P(B_1)+0.84\cdot P(B_2)+0.93\cdot P(B_3)}="


"={0.84\cdot P(B_2)\over 0.75\cdot P(B_2)+0.84\cdot P(B_2)+0.93\cdot P(B_2)}="

"={0.84\over 0.75+0.84+0.93}={28 \over 84}"



"P(B_3|A)={0.93\cdot P(B_3)\over 0.75\cdot P(B_1)+0.84\cdot P(B_2)+0.93\cdot P(B_3)}="


"={0.93\cdot P(B_3)\over 0.75\cdot P(B_3)+0.84\cdot P(B_3)+0.93\cdot P(B_3)}="

"={0.93\over 0.75+0.84+0.93}={31 \over 84}"